# Graham’s rearrangement for dihedral groups

### v1

@@ 4 Ordering the Set $E$
    If $(\nu_{k}+\pi_{1}(S_{o}^{k}),1)\not\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\min\{\pi_{1}(x):x\in S_{o}^{k}\mbox{ and }(\nu_{k}+\pi_{1}(x),1)\not\in Y_{i}\}.$
Similarly, in the second case, we choose $x_{k}$ as follows:
  1. 1.
    If there is some $x^{*}=(z^{*},1)\in S_{e}^{k}$ such that $(\nu_{k}-z^{*},0)\notin\cup_{j}Y_{j}$, then choose $x_{k}$ to be this $x^{*}$ and say that the current step is a skip-step for $S_{e}$.
  2. 2.
    Now, consider the case where $(\nu_{k}-\pi_{1}(S_{e}^{k}),0)\subseteq\cup_{j}Y_{j}$. Let $i$ be maximal such that $(\nu_{k}-\pi_{1}(S_{e}^{k}),0)$ intersects $Y_{i}$, and say that the current step is an $i$-step for $S_{e}$. If $(\nu_{k}-\pi_{1}(S_{e}^{k}),0)\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\max\{\pi_{1}(x):x\in S_{e}^{k}\}.$
    If $(\nu_{k}-\pi_{1}(S_{e}^{k}),0)\not\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\max\{\pi_{1}(x):x\in S_{e}^{k}\mbox{ and }(\nu_{k}-\pi_{1}(x),0)\not\in Y_{i}\}.$
In the third case, we choose $x_{k}$ as follows:
  1. 1.
    If there is some $x^{*}=(z^{*},0)\in N^{k}$ such that $(\nu_{k}+z^{*},1)\notin\cup_{j}Y_{j}$, then choose $x_{k}$ to be this $x^{*}$ and say that the current step is a skip-step for $N$.
  2. 2.
    Now, consider the case where $(\nu_{k}+\pi_{1}(N^{k}),1)\subseteq\cup_{j}Y_{j}$. Let $i$ be minimal such that $(\nu_{k}+\pi_{1}(N^{k}),1)$ intersects $Y_{i}$, and say that the current step is an $i$-step for $N$. If $(\nu_{k}+\pi_{1}(N^{k}),1)\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\min\{\pi_{1}(x):x\in N^{k}\}.$
    If $(\nu_{k}+\pi_{1}(N^{k}),1)\not\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\min\{\pi_{1}(x):x\in N^{k}\mbox{ and }(\nu_{k}+\pi_{1}(x),1)\not\in Y_{i}\}.$
Finally, in the fourth case, we choose $x_{k}$ as follows:
  1. 1.
    If there is some $x^{*}=(z^{*},0)\in P^{k}$ such that $(\nu_{k}-z^{*},0)\notin\cup_{j}Y_{j}$, then choose $x_{k}$ to be this $x^{*}$ and say that the current step is a skip-step for $P$.
  2. 2.
    Now, consider the case where $(\nu_{k}-\pi_{1}(P^{k}),0)\subseteq\cup_{j}Y_{j}$. Let $i$ be maximal such that $(\nu_{k}-\pi_{1}(P^{k}),0)$ intersects $Y_{i}$, and say that the current step is an $i$-step for $P$. If $(\nu_{k}-\pi_{1}(P^{k}),0)\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\max\{\pi_{1}(x):x\in P^{k}\}.$
    If $(\nu_{k}-\pi_{1}(P^{k}),0)\not\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\max\{\pi_{1}(x):x\in P^{k}\mbox{ and }(\nu_{k}-\pi_{1}(x),0)\not\in Y_{i}\}.$
@@
$|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leqslant 4\inf_{L\in\mathbb{N}}\left(\frac{|Y_{j}|}{L}+L+2\sum_{i=1}^{j-1}|Y_{i}|\right)$
for all $1\leqslant j\leqslant m$.
We split the set $\delta\cdot\operatorname{IS}(\mathbf{x})$ into the following ones:
$\delta\cdot\operatorname{IS}^{o}(\mathbf{x})=\{(\nu_{\ell},0):\ell\in[|N|+|P|+2,|N|+|P|+|S|]\mbox{ and }\ell-(|N|+|P|+|S|)\mbox{ is odd}\},$
$\delta\cdot\operatorname{IS}^{e}(\mathbf{x})=\{(\nu_{\ell},1):\ell\in[|N|+|P|+2,|N|+|P|+|S|]\mbox{ and }\ell-(|N|+|P|+|S|)\mbox{ is even}\}$
@@
$\delta\cdot\operatorname{IS}^{p}(\mathbf{x})=\{(\nu_{\ell},0):\ell\in[0,|P|]\},$
$\delta\cdot\operatorname{IS}^{n}(\mathbf{x})=\{(\nu_{\ell},1):\ell\in[|P|+2,|N|+|P|+1]\}.$
We will prove the statement of Equation (4) (without the factor four) only for $|\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}|$ since the argument for $|\delta\cdot\operatorname{IS}^{e}(\mathbf{x})\cap Y_{j}|$ is essentially identical and that of $|\delta\cdot\operatorname{IS}^{p}(\mathbf{x})\cap Y_{j}|$ and $|\delta\cdot\operatorname{IS}^{n}(\mathbf{x})\cap Y_{j}|$ is identical to that of [4].
Given such value of $\ell$, $(\nu_{\ell},0)$ lie in $Y_{j}$ only when the choice of $x_{\ell+1}$ is a $j$-step or an $i$-step (for $S_{e}$) for some $i<j$. We will estimate these two contributions separately. Note that skip-steps and $i$-steps for $i>j$ never contribute.
We first consider the contribution of $j$-steps. Suppose that the choice of $x_{\ell+1}=(z_{\ell+1},1)$ is a $j$-step for $S_{e}$ and $(\nu_{\ell},0)=(\nu_{\ell+1}-z_{\ell+1},0)\in Y_{j}$. Then we must have
$(\nu_{\ell+1}-\pi_{1}(S_{e}^{\ell+1}),0)\subseteq Y_{j}.$
@@
Due to the monotonicity of $\nu_{|P|+|N|+1+2t+1}$, setting $\ell=|P|+|N|+1+2t+1$, it follows that the $t-1$ elements of $(\nu_{\ell+1}-\pi_{1}(S_{e}^{\ell+1}),0)\setminus\{(\nu_{\ell},0)\}$ can never appear in $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})$. In particular, from such $j$-steps with $t\geqslant L+1$ we obtain at most $|Y_{j}|/L$ elements of $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}$. From $j$-steps with $t\leqslant L$ we trivially obtain at most $L$ elements of $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}$.
@@
We now consider the contribution of $i$-steps for $S_{e}$ with $i<j$. We will bound this contribution by the total number of $i$-steps for $S_{e}$ with $i<j$. We claim that the number of $i$-steps is at most $2|Y_{i}|$ for each $i$. For each $i$-step $\ell+1$, let $y(\ell)$ denote the smallest element of $\pi_{1}((\nu_{\ell+1}-\pi_{1}(S_{e}^{\ell+1}),0)\cap Y_{i})$. It suffices to show that each $y\in\pi_{1}(Y_{i})$ appears as $y(\ell)$ for at most $2$ different $i$-steps (for $S_{e}$).
If $y(\ell)$ is not the smallest element of $\nu_{\ell+1}-\pi_{1}(S_{e}^{\ell+1})$, then it is distinct from $y(\ell^{\prime})$ for all odd $\ell^{\prime}<\ell$ since
$y(\ell)>\nu_{\ell+1}-z_{\ell+1}=\nu_{\ell}$
and, given $z^{\prime}\in\pi_{1}(S_{e}^{\ell-1})$,
$\nu_{\ell}\geq\nu_{\ell}+(z_{\ell}-z^{\prime})=\nu_{\ell-1}-z^{\prime}\geq y_{\ell-2}.$
@@
Finally, set $\ell=|P|+|N|+1+2t+1$, note that the smallest element of $\nu_{\ell+1}-\pi_{1}(S_{e}^{\ell+1})$ is strictly increasing as $t$ increases. It follows that $y(\ell)$ is the smallest element of $\nu_{\ell+1}-\pi_{1}(S_{e}^{\ell+1})$ for at most one value of $\ell$. It suffices to show that each $y\in\pi_{1}(Y_{i})$ appears as $y(\ell)$ for at most $2$ different $i$-steps (for $S_{e}$), $\ell$. Summing up, we obtain that
@@
$|\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}|\leqslant\inf_{L\in\mathbb{N}}\left(\frac{|Y_{j}|}{L}+L+2\sum_{i=1}^{j-1}|Y_{i}|\right)$
@@
for all $1\leqslant j\leqslant m$. Since such an inequality holds also for $\delta\cdot\operatorname{IS}^{e}(\mathbf{x})$, $\delta\cdot\operatorname{IS}^{p}(\mathbf{x})$, and $\delta\cdot\operatorname{IS}^{n}(\mathbf{x})$ we get Equation (4). ∎
5 Ordering the Dissociated Sets
@@
Now the proof proceeds by ordering the dissociated sets, which are each constantly-one or constantly-zero in their $\mathbb{F}_{2}$ components, via a probabilistic procedure inspired by that of [4]. In particular, Lemma 5.1 of [4] allows us to split a dissociated set $D$ that is constantly-zero in its $\mathbb{F}_{2}$ components into uniform random partitions $D^{(1)},D^{(2)},D^{(3)}$, and $D^{(4)}$. Consequently, we need only consider the case of dissociated sets $D$ that are constantly-one in their $\mathbb{F}_{2}$ components in the following lemma. Note that we define $g=\Omega(f)$ if and only if $f\geq 0$ and $g\geq Cf$ for some constant $C>0$, and $g=O(f)$ if and only if $f\geq 0$ and $g\leq Cf$ for some constant $C>0$. Moreover, we define $g=\Theta(f)$ if and only if $g=O(f)$ and $g=\Omega(f)$.
@@
Let $D\subseteq Dih_{p}$ be a dissociated set such that $|D|$ is divisible by 8. We split $D$ into $D^{o}=\{d_{1}^{o},d_{2}^{o}\dots,d_{h}^{o}\}$ and $D^{e}=\{d_{1}^{e},d_{2}^{e}\dots,d_{h}^{e}\}$ so that $h=|D^{o}|=|D^{e}|=\frac{1}{2}|D|$. Then we let $D^{o}=D^{o,(1)}\cup D^{o,(2)}\cup D^{o,(3)}\cup D^{o,(4)}$ and $D^{e}=D^{e,(1)}\cup D^{e,(2)}\cup D^{e,(3)}\cup D^{e,(4)}$ be uniformly random partitions of $D^{o}$ and $D^{e}$ where each partition has the same cardinality. Then for every nonempty proper interval $I\subseteq\{1,2,3,4\}$ and each $x\in Dih_{p}$, we have
@@
$\prod_{\begin{subarray}{c}d^{o}\in D^{o,(i)}\cup D^{o,(j)}\cup D^{o,(k)}\\ d^{e}\in D^{e,(i)}\cup D^{e,(j)}\cup D^{e,(k)}\end{subarray}}d^{o}\cdot d^{e}$
@@
As in [4], we reorder the dissociated sets as follows. For each dissociated set $D_{j}$ that is constantly-one in its $\mathbb{F}_{2}$ (i.e. $j\in L_{1}$) components, we randomly partition $D_{j}=\cup_{i=1}^{4}D_{j}^{(i)}$ into four sets of equal size, where each $D_{j}^{(i)}=D_{j}^{o,(i)}\cup D_{j}^{e,(i)}$ as described in Lemma 5.1. For each dissociated set $D_{j}$ that is constantly-zero in its $\mathbb{F}_{2}$ components, we operate analogously to Lemma 5.1 of [4], therefore we partition uniformly at random $D_{j}=\cup_{i=1}^{4}D_{j}^{(i)}$ into four sets of equal size. All these partitions are performed independently. Then we order these new dissociated sets as follows
$D_{1}^{(1)},D_{1}^{(2)},D_{2}^{(1)},D_{2}^{(2)},\ldots,D_{s}^{(1)},D_{s}^{(2)},D_{1}^{(3)},D_{1}^{(4)},D_{2}^{(3)},D_{2}^{(4)},\ldots,D_{s}^{(3)},D_{s}^{(4)}\,.$ (5)
For notational convenience, we rename this new sequence as $T_{1},T_{2},\ldots,T_{u}$ where $u=4s$, and we define
$\tau_{j}:=\prod_{\begin{subarray}{c}d^{o}\in T_{j}^{o}\\ d^{e}\in T_{j}^{e}\end{subarray}}d^{o}\cdot d^{e}\,,$
if $T_{j}$ is a dissociated set that is constantly-one in its $\mathbb{F}_{2}$ components, otherwise we define it as
$\tau_{j}:=\prod_{d\in T_{j}}d\,.$
Let us also define
$\prod_{\leq m}(T_{j}):=\left\{\prod_{d^{o}\in S^{o},d^{e}\in S^{e}}d^{o}\cdot d^{e}\mid S^{o}\subseteq T_{j}^{o},S^{e}\subseteq T_{j}^{e},|S^{o}|=\lceil k/2\rceil,|S^{e}|=\lfloor k/2\rfloor,k\leq m\right\}$ (6)
if $T_{j}$ is a dissociated set that is constantly-one in its $\mathbb{F}_{2}$ components, otherwise we define it as
$\prod_{\leq m}(T_{j}):=\left\{\prod_{d\in S}d\mid S\subseteq T_{j},|S|\leq m\right\}\,.$ (7)
Given a subset $S\subseteq Dih_{p}$, we define $(S)^{-1}$ to be the set $\{d^{-1}\mid d\in S\}$.
Recalling the definition given in [4] of Type I and Type II intervals, we say that a proper nonempty interval $I\subset[1,|A|]$ is of Type II if it contains between $K$ and $|T_{j}|-K$ elements of some $T_{j}$, where $K=c_{2}R^{1/3}$ for a sufficiently small positive constant $c_{2}$; otherwise we say that the interval is of Type I.
In the following lemma, we adapt Lemma 5.3 of [4] to the dihedral case.
@@
Let $p$ be large enough and let $c>0$ and $1\leq s\leq e^{c(\log p)^{1/4}}$. Let $D_{1},D_{2},\ldots,D_{s}\subseteq Dih_{p}$ be dissociated sets, each of size $\Theta(R)$. Let $\mathbf{x}$ be a sequence over $Dih_{p}$ of length at most $e^{c(\log p)^{1/4}}$, and assume that $\delta,\mathbf{x}$ is a sequencing. If the sequence $T_{1},\ldots,T_{u}$ of dissociated sets is chosen randomly as described previously, then we have with positive probability:

### v2

@@ 4. Ordering the Set $E$ and the Dissociated Sets
    If $\nu_{k}+S_{o}^{k}\not\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\min\{\pi_{1}(x):x\in S_{o}^{k}\mbox{ and }\nu_{k}+x\not\in Y_{i}\}.$
In the second, third, and fourth cases, we choose $x_{k}\in S^{k}_{e},N^{k},P^{k}$, respectively, by the same method as in case 1 where $x_{k}$ is chosen as a minimum in the case of $N^{k}$ and as a maximum otherwise.
@@
$|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leq 4\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)+|H|$
for all $1\leq j\leq m$.
We split the set $\delta\cdot\operatorname{IS}(\mathbf{x})$ into the following subsets:
$\delta\cdot\operatorname{IS}^{o}(\mathbf{x})=\{\nu_{\ell}:\ell\in[|N|+|P|+2,|N|+|P|+|S|]\mbox{ and }\ell-(|N|+|P|+|S|)\mbox{ is odd}\},$
$\delta\cdot\operatorname{IS}^{e}(\mathbf{x})=\{\nu_{\ell}:\ell\in[|N|+|P|+2,|N|+|P|+|S|]\mbox{ and }\ell-(|N|+|P|+|S|)\mbox{ is even}\}$
@@
$\delta\cdot\operatorname{IS}^{p}(\mathbf{x})=\{\nu_{\ell}:\ell\in[0,|P|]\},$
$\delta\cdot\operatorname{IS}^{n}(\mathbf{x})=\{\nu_{\ell}:\ell\in[|P|+2,|N|+|P|+1]\}.$
Equation (9) follows from $|\delta\cdot\operatorname{IS}(\mathbf{z})|\leq|H|$ and by showing that $|\delta\cdot\operatorname{IS}^{p}(\mathbf{x})\cap Y_{j}|$, $|\delta\cdot\operatorname{IS}^{n}(\mathbf{x})\cap Y_{j}|$, $|\delta\cdot\operatorname{IS}^{e}(\mathbf{x})\cap Y_{j}|$, $|\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}|$ are upper bounded by $\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)$. We will prove the statement of Equation (9) (without the factor of four) only for $|\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}|$ since the argument for $|\delta\cdot\operatorname{IS}^{e}(\mathbf{x})\cap Y_{j}|$ is essentially identical and that of $|\delta\cdot\operatorname{IS}^{p}(\mathbf{x})\cap Y_{j}|$ and $|\delta\cdot\operatorname{IS}^{n}(\mathbf{x})\cap Y_{j}|$ is identical to that of [4].
Given some value of $\ell$, $\nu_{\ell}$ lie in $Y_{j}$ only when the choice of $x_{\ell+1}$ is a $j$-step or an $i$-step (for $S_{e}$) for some $i<j$. We will estimate these two contributions separately. Note that skip-steps and $i$-steps for $i>j$ never contribute.
We first consider the contribution of $j$-steps. Suppose that the choice of $x_{\ell+1}=(z_{\ell+1},a_{\ell+1})$ is a $j$-step for $S_{e}$ and $\nu_{\ell}=\nu_{\ell+1}-x_{\ell+1}\in Y_{j}$. Then we must have
$\nu_{\ell+1}-S_{e}^{\ell+1}\subseteq Y_{j}.$
@@
Due to the monotonicity of $\pi_{1}(\nu_{|P|+|N|+1+2t+1})$, setting $\ell=|P|+|N|+1+2t+1$, it follows that the (t1)/|H|(t-1)/|H| elements of $\nu_{\ell+1}-S_{e}^{\ell+1}\setminus\{\nu_{\ell}\}$ can never appear in $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})$. In particular, from such $j$-steps with $t\geq L+1$ we obtain at most $|H||Y_{j}|/L$ elements of $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}$. From $j$-steps with $t\leq L$ we trivially obtain at most $L$ elements of $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}$.
@@
We now consider the contribution of $i$-steps for $S_{e}$ with $i<j$. We will bound this contribution by the total number of $i$-steps for $S_{e}$ with $i<j$. We claim that the number of $i$-steps is at most $2|H||Y_{i}|$ for each $i$. For each $i$-step $\ell+1$, let $y(\ell)$ denote the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1}\cap Y_{i})$. It suffices to show that each $y\in\pi_{1}(Y_{i})$ appears as $y(\ell)$ for at most $2|H|$ different $i$-steps (for $S_{e}$).
If $y(\ell)$ is not the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1})$, then it is distinct from $y(\ell^{\prime})$ for all odd $\ell^{\prime}<\ell$ since
$y(\ell)>\pi_{1}(\nu_{\ell+1})-z_{\ell+1}=\pi_{1}(\nu_{\ell})$
and, given $z^{\prime}\in\pi_{1}(S_{e}^{\ell-1})$, by Equation (7) we have
$\pi_{1}(\nu_{\ell})\geq\pi_{1}(\nu_{\ell})+(z_{\ell}-z^{\prime})=\pi_{1}(\nu_{\ell-1})-z^{\prime}\geq y(\ell-2).$
@@
Finally, set $\ell=|P|+|N|+1+2t+1$ and note that the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1})$ is nondecreasing as $t$ increases. It follows that $y(\ell)$ is the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1})$ for at most $|H|$ values of $\ell$. Summing up, we obtain that
@@
$|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leq 4\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)+|H|$
@@
for all $1\leq j\leq m$. Since such an inequality holds also for $\delta\cdot\operatorname{IS}^{e}(\mathbf{x})$, $\delta\cdot\operatorname{IS}^{p}(\mathbf{x})$, and $\delta\cdot\operatorname{IS}^{n}(\mathbf{x})$, we get Equation (9). ∎
With essentially the same proof of Proposition 4.2 and thanks to Claims 4.2 and 4.3 of [4] we obtain the following proposition which shows that, for any family of subsets $Y_{1},\ldots,Y_{m},Y^{\prime}_{1},\ldots,Y^{\prime}_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$, we can order $E$ in the case of $S=\emptyset$ as follows:
Proposition 4.3.
Let $p$ be large enough. Consider $\delta=(z_{0},0)\in\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $z_{0}>0$ and subsets $Y_{1},\ldots,Y_{m},Y^{\prime}_{1},\ldots,Y^{\prime}_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$. Take a finite subset $E$ such that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ and consider $P,N,Z,S$ to be defined as above where $S=\emptyset$. Then there are orderings $\mathbf{p}$ of $P$, $\mathbf{n}$ of $N$, and $\mathbf{z}$ of $Z$ such that $\mathbf{z},\overline{\mathbf{p}},\delta,\mathbf{n}$ is a sequencing and
(10) $|\delta\cdot\operatorname{IS}(\mathbf{p})\cap Y_{j}|\leq\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)$
(11) $|\delta\cdot\operatorname{IS}(\mathbf{n})\cap Y^{\prime}_{j}|\leq\inf_{L\in\mathbb{N}}\left(\frac{|H||Y^{\prime}_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y^{\prime}_{i}|\right)$
for all $1\leq j\leq m$.
@@
Now the proof proceeds by ordering the dissociated sets via a probabilistic procedure inspired by that of [4]. In particular, Lemma 5.1 of [4] allows us to split a dissociated set $D$, where the $H$ component of each element of $D$ is in $H_{0}$, into uniform random partitions $D^{(1)},D^{(2)},D^{(3)}$, and $D^{(4)}$. Consequently, we need only consider the case of dissociated sets $D$ with $H$ components in $H_{1}$. Note that we use the notation $\Omega$ as defined in [4].
@@
Let $D\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ be a dissociated set with $H$ components in $H_{1}$ and where $|D|$ is divisible by $8$. We split $D$ into $D^{o}=\{d_{1}^{o},d_{2}^{o}\dots,d_{h}^{o}\}$ and $D^{e}=\{d_{1}^{e},d_{2}^{e}\dots,d_{h}^{e}\}$ where $h=|D^{o}|=|D^{e}|=\frac{1}{2}|D|$. Then we let $D^{o}=D^{o,(1)}\cup D^{o,(2)}\cup D^{o,(3)}\cup D^{o,(4)}$ and $D^{e}=D^{e,(1)}\cup D^{e,(2)}\cup D^{e,(3)}\cup D^{e,(4)}$ be uniformly random partitions of $D^{o}$ and $D^{e}$, where each partition has the same cardinality. Then for every nonempty proper interval $I\subseteq\{1,2,3,4\}$ and each $x\in\mathbb{Z}_{p}\rtimes_{\varphi}H$, we have
@@
$\prod_{\begin{subarray}{c}d^{o}\in D^{o,(i)}\cup D^{o,(j)}\cup D^{o,(k)}\\ d^{e}\in D^{e,(i)}\cup D^{e,(j)}\cup D^{e,(k)}\end{subarray}}d^{o}\cdot d^{e}.$
@@
As in [4], we reorder the dissociated sets as follows. For each dissociated set $D_{j}$ with $j\in L_{1}$, we randomly partition $D_{j}=\cup_{i=1}^{4}D_{j}^{(i)}$ into four sets of equal size, where each $D_{j}^{(i)}=D_{j}^{o,(i)}\cup D_{j}^{e,(i)}$ as described in Lemma 4.4. For each dissociated set $D_{j}$ with $j\in L_{0}$, we operate analogously to Lemma 5.1 of [4] in that we partition uniformly at random $D_{j}=\cup_{i=1}^{4}D_{j}^{(i)}$ into four sets of equal size. All these partitions are performed independently. Then we order these new dissociated sets as follows:
(12) $D_{1}^{(1)},D_{1}^{(2)},D_{2}^{(1)},D_{2}^{(2)},\ldots,D_{s}^{(1)},D_{s}^{(2)},D_{1}^{(3)},D_{1}^{(4)},D_{2}^{(3)},D_{2}^{(4)},\ldots,D_{s}^{(3)},D_{s}^{(4)}\,.$
For notational convenience, we rename this new sequence as $T_{1},T_{2},\ldots,T_{u}$ where $u=4s$, and we use the same notation $T_{j}^{o}$ and $T_{j}^{e}$ as used for the $D_{j}$s. We define $\tau_{j}:=\prod_{d^{o}\in T_{j}^{o},d^{e}\in T_{j}^{e}}d^{o}\cdot d^{e}$ if $T_{j}$ is a dissociated set associated to a $D_{k}$ with $k\in L_{1}$, otherwise we define it as $\tau_{j}:=\prod_{d\in T_{j}}d$. We use $\mathbf{t}_{i}$ to denote an ordering of $T_{i}$ that is chosen according to a suitable distribution. Let us also define
(13) $\prod_{\leq m}(T_{j}):=\left\{\prod_{d^{o}\in S^{o},d^{e}\in S^{e}}d^{o}\cdot d^{e}:S^{o}\subseteq T_{j}^{o},S^{e}\subseteq T_{j}^{e},|S^{o}|=\lceil k/2\rceil,|S^{e}|=\lfloor k/2\rfloor,k\leq m\right\}$
if $T_{j}$ is a dissociated set associated to a $D_{k}$ with $k\in L_{1}$. Otherwise, we define it as
(14) $\prod_{\leq m}(T_{j}):=\left\{\prod_{d\in S}d:S\subseteq T_{j},|S|\leq m\right\}\,.$
Given a subset $S$ of a finite group, we define $(S)^{-1}$ to be the set $\{d^{-1}:d\in S\}$.
Recalling the definition given in [4] of Type I and Type II intervals, we say that a proper nonempty interval $I\subset[1,|A|]$ is of Type II if it contains between $K$ and $|T_{j}|-K$ elements of some $T_{j}$, where $K=c_{2}R^{1/3}$ for a sufficiently small positive real $c_{2}$; otherwise we say that the interval is of Type I.
In the following lemma, we adapt Lemma 5.3 of [4] to the semidirect product case.
@@
Let $p$ be large enough and let $c>0$ and $1\leq s\leq e^{c(\log p)^{1/4}}$. Take a sequence of dissociated sets $D_{1},D_{2},\ldots,D_{s}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ each of size $\asymp R$ and a multiple of eight, where $H$ is abelian and $\varphi$ satisfies property $(\star)$. Let $\mathbf{x}_{1},\mathbf{x}_{2}$ be sequences over $\mathbb{Z}_{p}\rtimes_{\varphi}H$ of length at most $e^{c(\log p)^{1/4}}$, and assume that $\mathbf{x}_{1},\delta,\mathbf{x}_{2}$ is a sequencing. If the sequence $T_{1},\ldots,T_{u}$ of dissociated sets is chosen randomly as described previously, then we have with positive probability that